### v1

# WebDevJudge: Evaluating (M)LLMs as Critiques for Web Development Quality

Abstract
The paradigm of LLM-as-a-judge is emerging as a scalable and efficient alternative to human evaluation, demonstrating strong performance on well-defined tasks. However, its reliability in open-ended tasks with dynamic environments and complex interactions remains unexplored. To bridge the gap, we introduce WebDevJudge, a systematic benchmark for assessing LLM-as-a-judge performance in web development, with support for both non-interactive evaluation based on static observations and continuous interactive evaluation with a dynamic web environment. WebDevJudge comprises human preference labels over paired web implementations, annotated with structured and query-grounded rubrics to ensure high-quality ground truth. Using this benchmark, we comprehensively evaluate various evaluators, including LLMs, MLLMs, and agentic workflows. We systematically investigate the impact of different paradigms and guidance mechanisms. Our experiments reveal a significant gap between LLM judges and human experts. In-depth analysis indicates this gap stems from fundamental model limitations, including failures in recognizing functional equivalence, verifying task feasibility, and mitigating bias. Overall, WebDevJudge presents a significant challenge to LLM-as-a-judge, offering insights to guide future research toward developing more reliable and capable automated evaluators for complicated scenarios. Code and data are available at https://github.com/lcy2723/WebDevJudge.
1 Introduction
Evaluate, refine, then evaluate again and refine again—large language models (LLMs) have achieved remarkable success across various domains through this iterative cycle (Madaan et al., 2023; Chen et al., 2023; Shafayat et al., 2025; Li et al., 2025a). Conventional evaluation paradigms rely heavily on human assessment (Zhong et al., 2024; Starace et al., 2025), which, while meticulous, presents a critical bottleneck due to its high cost and low scalability. In response, the paradigm of LLM-as-a-judge has emerged as a promising alternative (Zheng et al., 2023; Dubois et al., 2024), offering a scalable and cost-effective solution for crucial development stages like verification (Lightman et al., 2024) and reward modeling (Yuan et al., 2025). With the advent of sophisticated language agents capable of planning, tool use, and collaboration (Yao et al., 2023; Qin et al., 2024a; Liang et al., 2024), the role of LLM-as-a-judge is rapidly expanding beyond basic, well-defined tasks to encompass challenging real-world problems (Zhuge et al., 2025; Bian et al., 2025). This progression is critical, as it paves the way for more capable automated evaluators, creating the possibility for language models to self-evolve in complex, real-world applications.
However, a fundamental question regarding the reliability of LLM-as-a-judge persists. Its effectiveness is well-established for static and basic tasks (Chen et al., 2024; Saha et al., 2025; Gou et al., 2025; Lù et al., 2025), but these successes share a critical commonality: they rely on static assessment of final outcomes. The reliability of LLM-as-a-judge in dynamic, open-ended domains involving complex interaction remains largely unexplored. Such contexts introduce significant challenges: dynamic evaluation requires continuous interaction with and comprehension of a changing environment (Li et al., 2023a; Paglieri et al., 2025), while their open-ended nature necessitates establishing feasible assessment standards (Li et al., 2025b). This gap between the expanding scope of automated judges and the lack of rigorous validation in complex, interactive settings highlights an urgent need for a new meta-evaluation benchmark. Such a benchmark is essential to assess, interpret, and ultimately enhance the reliability of LLM-based judges as we approach increasingly autonomous AI systems.
Consequently, in this work, we introduce WebDevJudge, a meta-evaluation benchmark for assessing LLM-as-a-judge on a representative complex and interactive task using the context of web development. Web development offers an ideal testbed for complex dynamic evaluation, as it inherently has interaction requirements, whose assessment depends not just on static code but on real-time interaction. The task is also intrinsically open-ended, lacking a single absolute answer. To establish a high-quality ground truth, we introduce a structured annotation methodology using query-grounded rubric trees, which decomposes high-level requirements into a verifiable hierarchy of fine-grained criteria. This rigorous protocol combining rubric with human judges yields an inter-annotator agreement of 89.7%, which substantially exceeds the 63% reported for MT-Bench (Zheng et al., 2023), confirming the reliability of the preference labels. Departing from traditional benchmarks focused on static text (Tan et al., 2025), WebDevJudge supports both static and interactive evaluation by offering multifaceted representations of each web implementation, including its source code, screenshot of the rendered webpage, and a fully interactive environment for dynamic assessment, as shown in the right part of Figure 1. Evaluator performance is measured by agreement with expert human preferences on paired web implementations—a methodology widely adopted for nuanced, open-ended tasks where the absolute answer is deficient (Bai et al., 2022; Lambert et al., 2025).
Using WebDevJudge, we conduct a comprehensive evaluation of a wide array of judges, including LLMs, MLLMs, and agentic workflows, under various paradigms and guidance mechanisms. Our experiments reveal that a significant capabilities gap persists between the most advanced models and human experts, with a performance discrepancy exceeding 15%. Notably, different guidance strategies provide only marginal improvements in the pairwise comparison setting, suggesting that preference prediction through comparative assessment represents an internalized capability in models (Hua et al., 2025; Yu et al., 2025). Furthermore, while agentic workflows appear well-suited for interactive task evaluation, they fail to outperform vanilla models due to error accumulation across planning and execution stages (Pan et al., 2025). Through detailed error analysis and case studies, we systematically investigate the failure modes of automated evaluators. As part of this analysis, we construct WebDevJudge-Unit, a diagnostic dataset specifically designed to evaluate feasibility verification capabilities of different types of LLM-as-a-judge. Our investigation reveals two fundamental performance bottlenecks: (1) a persistent inability to recognize functional equivalence between diverse implementations that achieve the same objectives through different approaches or terminologies, such as implementations using the same title element with only variations in text, and (2) systematic weaknesses in feasibility verification, where static assessment suffers from low precision due to static code analysis limitations while interactive agents exhibit low recall stemming from their own operational constraints. These compounding errors ultimately undermine evaluator performance, pointing toward fundamental research directions for developing truly reliable automated evaluators in complex, interactive domains. Our main contributions are as follows:
  • We construct WebDevJudge, a meta-evaluation benchmark that supports both static and interactive assessment of web development quality with high-quality preference labels.
  • Comprehensive empirical evaluation of (M)LLMs and agentic workflow reveals that current LLM-as-a-judge approaches still fall significantly short of human-level reliability.
  • Detailed error analysis identifies the systematic weaknesses of LLM-as-a-judge, providing critical insights for developing more reliable automated evaluators.
2 Related work
LLM-as-a-Judge
The paradigm of LLM-as-a-Judge leverages powerful large language models to simulate human-like assessment, enabling scalable and cost-efficient evaluation (Gu et al., 2025). This approach has been widely adopted across various domains, including question generation and answering (Bai et al., 2023), data filtering (Li et al., 2024; Xu et al., 2025), and trajectory evaluation (Pan et al., 2024; Xue et al., 2025). Typical implementations employ a strong LLM as an evaluator, which compares or scores candidate responses against predefined criteria. Common comparison methods include pairwise comparison and single-answer grading Zheng et al. (2023). As application scenarios broaden, the conventional LLM-as-a-Judge approach is evolving into Agent-as-a-Judge (Zhuge et al., 2025), wherein LLM-based agents are equipped with tool-using and collaboration capabilities. Gou et al. (2025) propose a judge agent with extractor and verifier to evaluate deep research tasks with rubric trees, while Zhuge et al. (2025) design a modular agentic framework to evaluate agentic systems. Despite its efficiency, this method faces several challenges. Position bias (Wang et al., 2024) and verbosity (Ye et al., 2025) preference may skew results, and the creation of detailed, human-written rubrics is both labor-intensive and difficult to scale (Starace et al., 2025). Additionally, typical assessments may overemphasize final outcomes, limiting their applicability in open-ended interactive tasks (Zhang et al., 2025a). To evaluate the judge capability of LLMs, we introduce a benchmark on open-ended web development scenarios. It is designed to assess a wide range of LLM-based evaluators, with the dual purpose of benchmarking their performance and exposing fundamental shortcomings in existing automated evaluation approaches.
Meta Evaluation
Meta-evaluation accesses the effectiveness of automated evaluation by measuring its correlation with human or ground truth. Existing benchmarks primarily focus on two aspects: (1) alignment with preference label, often using pairwise comparisons (Zheng et al., 2023; Li et al., 2023b), and (2) accuracy in identifying correct task outcomes, particularly in reasoning (Luo et al., 2023; He et al., 2025) and agentic tasks (Lù et al., 2025). Prior works such as MT-bench (Zheng et al., 2023) and LLMEval (Zhang et al., 2023) measure how LLM judges mirror human preferences in multi-turn conversational and instruction following tasks. However, these benchmarks may be constrained by human subjectivity and position bias (Wang et al., 2024). Alternatives like LLMBar (Zeng et al., 2024) and JudgeBench (Tan et al., 2025) test judges on re-annotated instruction adherence or verifiable reasoning discrimination. These benchmarks mainly focus on text-based tasks without complex environments. In the context of interaction, AgentRewardBench (Lù et al., 2025) uses expert-annotated result of pre-scripted trajectories to benchmark how well LLM evaluators score agent performance. ArtifactsBench (Zhang et al., 2025a) evaluates dynamic visual effects. However, it lacks a critical component of assessing environmental changes that are driven by real-time user input. Our work, WebDevJudge, introduces a meta-evaluation benchmark that assesses judges on dynamic, real-world web development tasks, emphasizing continuous interaction with live web environments, offering insights in complex interactive settings.
Figure 1: Overview of WebDevJudge. Left: Data Collection with filtering. Center: Annotation with Rubric. Right: Evaluate various evaluators based on different evaluation paradigms.
3 WebDevJudge Benchmark
WebDevJudge serve as a meta-evaluation benchmark designed to assess whether LLM-as-a-judge can effectively approximate human preference judgments in web development tasks. We frame the task as a preference evaluation problem. Each instance is represented as a quadruple $(Q,W_{\text{a}},W_{\text{b}},l_{\text{p}})$, where $Q$ denotes a web development query (e.g., “build a book review page”), $W_{\text{a}}$ and $W_{\text{b}}$ represent web implementations from two distinct models $a$ and $b$, and $l_{\text{p}}$ is the annotated preference label between the two outputs. The objective is to predict the preference label—i.e., whether $a$ wins, $b$ wins, or the result is a tie—which is widely used for open-ended tasks (Zeng et al., 2024; Lambert et al., 2025). The observations of $W_{\text{a}}$ and $W_{\text{b}}$ may take various forms across experimental settings, such as code, screenshots, or interaction trajectories.
3.1 Data Filtering
We collect data from the webdev-arena-preference-10k dataset (Vichare et al., 2025; Chiang et al., 2024), which comprises 10,501 user queries with paired code outputs from 2 models and user-provided preferences. We apply a two-stage filtering process to enhance data quality:
Query-based filtering
We first exclude extremely short queries and verbatim duplicates. The remaining queries are further filtered via an LLM according to the following criteria: (1) Safety: exclusion of harmful or offensive content; (2) Clarity: the query must be unambiguous and interpretable; (3) Feasibility: the query should be realistically implementable as a web application, based on its purpose and required level of interaction.
Environment-based filtering
We deploy each web implementation in a unified execution environment. Instances that fail to deploy correctly or require niche dependencies are discarded. To ensure validity, we capture initial screenshots of each rendered webpage and use a multi-modal LLM to filter out invalid cases (e.g., blank pages or intrinsic errors).
After applying these filters, we retain 1,713 high-quality instances. Further details regarding the data construction process are provided in Appendix A.
Table 1: Agreement rates with and without the rubric-based annotation. The ‘without rubric’ part shows agreements between: (1) annotators and (2) annotators and the original labels. The ‘with rubric’ part shows inter-annotator agreements under rubrics from different sources.
Setting Without rubric With rubric
inter-annotator annotator-origin Human-written LLM-generated
w/ tie 65.0 53.0 92.0 90.0
w/o tie 90.0 77.9 95.5 96.9
3.2 Annotation via Rubric tree
Consistent with prior work (Zheng et al., 2023; Tan et al., 2025), we note that raw preference labels may reflect subjective bias rather than objective quality. To quantify this issue, we sample 100 instances and have them re-annotated by two expert annotators based on fully deployed web instances. As shown in Table 1, both inter-annotator agreement and agreement with original labels are initially low, highlighting the need for a more structured and standard annotation protocol.
To address this, we introduce rubric tree—a structured evaluation framework commonly used in complex assessment scenarios (Starace et al., 2025; Gou et al., 2025). The tree is query-based and scalable by design, organized along three core dimensions: intention, static quality, and dynamic behavior (see Figure 1). Each leaf node corresponds to a binary test, whose outcomes are aggregated hierarchically to parent nodes. This allows for both a holistic score at the root and fine-grained diagnostic insight via leaf-level pass rates.
We validate the effectiveness of the rubric tree by manually constructing trees for 50 instances and asking two annotators to evaluate these using the rubric as a reference. As shown in Table 1, rubric-guided annotation significantly improves agreement rates. To scale this process, we employ few-shot LLM generation to automatically produce rubric trees. A third annotator labels the same instances using these generated rubrics, achieving high agreement with human-written rubrics—confirming the utility of LLM-generated rubrics. We then use generated rubric trees to annotate the remaining data via two expert annotators with software engineering backgrounds. During annotation, we perform manual inspections to exclude incompatible and harmful cases. Final inter-annotator agreement reaches 89.7% (with ties) and 94.0% (without ties), demonstrating high consistency.
Table 2: Categories and their respective sub-categories of queries in WebDevJudge.
Category Sub-categories Total % Digital Design Website Design UI Design 34.0 Game and App Development Game Development App Development App Design 34.4 Web and Specialized Technologies Clone Development Web Development Simulations AI Applications Multilingual Queries Digital Tools Creative Humor 31.6
Figure 2: The distribution of sub-category across WebDevJudge.
3.3 Data Statistics
The final benchmark consists of 654 instances, with preference distributions as follows: 258 for $a$, 258 for $b$, and 138 ties. To characterize the query domain coverage, we conduct a topic analysis based on the original topics webdev-arena-preference-10k dataset (Vichare et al., 2025). Fine-grained topics are grouped into three broad categories: Digital Design, Game and App Development, and Web and Specialized Technologies based on their shared characteristics and application contexts of representative examples. Table 2 and Figure 2 provide detailed category distributions and visualizations.
4 Experiments
We conduct comprehensive experiments and evaluate different evaluation settings on WebDevJudge, focusing on the following research questions: (1) Whether LLM-as-a-judge can be an alternative to human preference in open-ended complex tasks like web development? (2) How do different settings and strategies affect the agreement rate performance?
4.1 Evaluator Performance on WebDevJudge
Setup
Following (Zheng et al., 2023; Xie et al., 2025), we evaluate the evaluators under 2 distinct paradigms:
  • Pairwise comparison: This approach directly compares two responses, leveraging relative judgments to infer preference, an intuitive and natural fit for preference prediction.
  • Single answer grading: This approach employs LLMs to assign scores or labels to individual responses, with preferences derived by comparing scores or pass rates.
To enable structured and domain-grounded assessment, we design a multi-dimensional Likert scale inspired by prior work (Lan et al., 2024; Zhang et al., 2025a) and incorporate principles from international software testing standards (ISO/IEC/IEEE, 2022). Our criteria comprises four dimensions, Functionality, UI Quality, Code Quality, and Interactivity, each with several sub-criteria rated on a 5-point Likert scale (1: lowest, 5: highest). Detailed criteria are provided in Appendix B.
Implementation
We evaluate two types of evaluators: (1) Vanilla (M)LLMs: we test models from OpenAI (OpenAI, 2025; 2024), Anthropic (Anthropic, 2025b; a), Google (Team, 2025a), Qwen (Team, 2025e; d), DeepSeek (DeepSeek-AI, 2024; 2025), Moonshot (Team, 2025c), and ZAI (Team, 2025b). For single answer grading, we provide the query, code, and evaluation criteria; we further include an initial screenshot of the rendered webpage for MLLMs. The model is prompted to output scores for each sub-criterion. For pairwise comparison, we supply the query, code of $W_{\text{a}}$ and $W_{\text{b}}$, evaluation criteria, and screenshots of both rendered webpages (if applicable). The evaluator scores both $W_{\text{a}}$ and $W_{\text{b}}$ per sub-criterion, with explicit instructions to ignore position bias and remain objective. Preference is determined by comparing total scores. (2) Agentic Workflow: similar to Bian et al. (2025), we model evaluation as a multi-stage pipeline with a planner, an executor, and a summarizer:
$\text{Query}\xrightarrow{\text{{Planner}}}\text{Plan with test cases}\xrightarrow{\text{{Executor}}}\text{Results}\xrightarrow{\text{{Summarizer}}}\text{Judge}$ (1)
The planner generates a verifiable evaluation plan conditioned on the query. The executor runs test cases, and the summarizer synthesizes results into a judgment. We use the generated rubric tree as the evaluation plan and reference for summarization. Preferences are inferred by comparing summarized outcomes. In implementation, we utilize UI-TARS-1.5 (Seed, 2025), one of the state-of-the-art GUI agents, as the executor, to ensure the reliability of the evaluation process.
Table 3: Agreement Rate (%) of different evaluators under different evaluation paradigms. The best average performance of the whole dataset is highlighted in bold and the second best is underlined.
Model/Method Digital Design Game & App Web & Special Average Single Pair Single Pair Single Pair Single Pair Vanilla Non-reasoning Models  GPT-4.1 54.95 65.77 61.33 67.56 60.39 64.73 58.87 66.06  GPT-4o 51.80 64.41 56.00 64.44 60.39 64.25 55.96 64.37  Qwen-2.5-VL-72B-Inst. 46.40 64.86 48.44 62.67 52.17 63.29 48.93 63.61  Gemini-2.5-flash-lite 48.20 62.16 48.89 57.78 47.34 62.80 48.17 60.86 DeepSeek-V3-0324 58.11 63.96 54.22 61.33 59.90 65.70 57.34 63.61 Kimi-K2-Inst. 62.16 64.41 56.00 61.78 52.17 65.22 56.88 63.76 Qwen3-235B-A22B-Inst. 53.60 63.51 59.56 64.00 58.45 61.35 57.19 63.00 Qwen3-30B-A3B-Inst. 54.50 64.41 58.67 60.44 55.07 65.22 56.12 63.30 Qwen2.5-32B-Inst. 48.65 61.26 53.33 64.89 51.69 61.35 51.22 62.54 Gemma-3-27B-it 44.14 59.01 52.00 54.67 45.41 60.39 47.25 57.95 Qwen2.5-14B-Inst. 48.20 56.31 44.00 51.56 38.65 52.17 43.73 53.36 Reasoning Models  Claude-4-Sonnet 50.90 67.12 58.67 68.44 57.00 62.32 55.50 66.06  Claude-3.7-Sonnet 65.77 67.12 59.11 61.33 57.49 67.15 60.86 65.14  Gemini-2.5-pro 51.35 61.26 50.22 67.11 43.48 57.97 48.47 62.23 GLM-4.5 57.66 66.22 52.44 64.44 50.72 60.87 53.67 63.91 DeepSeek-R1-0528 50.00 64.41 56.00 63.11 49.76 61.84 51.99 63.15 Qwen3-30B-A3B-Think. 54.05 59.91 53.33 59.11 52.17 60.39 53.21 59.79 Agentic Intention rubrics 37.39 - 32.00 - 47.34 - 38.69 - Static rubrics 40.09 - 38.67 - 43.48 - 40.67 - Dynamic rubrics 52.70 - 55.56 - 50.72 - 53.06 - Combined rubrics 54.50 - 56.44 - 57.49 - 56.12 - Human Pairwise comparison - 85.78 - 85.02 - 83.78 - 84.82
Result
The main experimental results are summarized in Table 3. Our analysis yields several key observations on the performance of LLM-based evaluators in web development tasks.
LLM-as-a-judge falls short of human-level reliability on complex evaluation tasks. The primary finding is that no current model achieves a sufficient level of agreement with expert judgments. The top-performing evaluator, Claude-4-Sonnet under the pairwise paradigm, attains an agreement rate of only 66.06%. This substantial gap underscores the inherent complexity of web development evaluation, which demands a holistic assessment of functionality and aesthetics. We also observe a clear performance ceiling: while smaller models exhibit scaling effects, larger and more capable models show diminishing returns, consistently plateauing below the mid-70% agreement rate.
Pairwise comparison is a far more effective paradigm for preference evaluation. Across the board, the pairwise paradigm yields an average improvement of over 8.0% in agreement rate compared to single-answer grading. Relative judgment helps models focus on discriminative features between two candidates, reducing the need for absolute quality calibration. In contrast, single-answer grading not only performs worse but also leads to inconsistent model rankings—some models deviate from their expected benchmark ordering (Zhang et al., 2025b), indicating that this paradigm is less suitable for open-ended, multi-faceted tasks. This may be due to the cognitive load of applying multi-point Likert scales consistently (Ouwehand et al., 2021), which demands a calibrated internal standard of quality that is difficult for both humans and LLMs to maintain. Simplified evaluation mechanisms, such as binary checklists, may be more effective; we explore this further in Section 4.2.
Agentic workflow suffers from compounding errors. We evaluated the agentic workflow using rubrics from different aspects, as well as a comprehensive score integrating all three. The results show that it performs significantly better on the Dynamic aspect, which involves strong interactive properties, compared to others with weaker interactivity. Counter-intuitively, the agentic workflow fails to outperform the vanilla models. This appears to result from error accumulation across its multi-stage process. We identify two primary failure modes:
  • Brittle Planning: The Planner struggles with the ambiguity of user queries that are often not expressed with expert-level precision. This leads to the generation of evaluation plans that are either too generic to be discriminative or overly specific, causing failures due to minor implementation variations.
  • Faulty Execution: The Executor agent’s ability to navigate the web and verify task states remains unreliable. It may misinterprets outcomes, injecting noise into the evaluation process. This unreliability is investigated more deeply in Section 4.3.
Errors compound across the planner-executor-summarizer pipeline, reducing the reliability of the final judgment compared to an end-to-end evaluation by a single model (Chen et al., 2025b).
4.2 Different Evaluation Guides and Observation Forms
To further diagnose factors affecting evaluator performance, we conduct controlled experiments focusing on two key variables: the structure of evaluation guidance and the form of observation.
Figure 3: Agreement rates of LLM evaluators under various guidance protocols. We compare pairwise and single-answer paradigms using direct judgment, Likert scales, and structured rubrics.
Evaluation Guidance
We investigate how different evaluation protocols influence judgment accuracy, comparing three widely-used forms of guidance (Zheng et al., 2023; Gou et al., 2025): (1) Direct: The evaluator provides a preference without explicit criteria. (2) Likert Scale: As used in Section 4.1, the evaluator scores outputs along predefined dimensions using a multi-point scale. (3) Rubric: The evaluator assesses outputs using the rubric trees from our annotation process, with final scores computed as a weighted aggregation of binary leaf-node pass rates across three core dimensions (intention, static, and dynamic).
Results in Figure 3 reveal a noteworthy finding: under the pairwise evaluation, the Direct setting achieves agreement rates comparable to guidance-based methods. This phenomenon aligns with observations in instruction following tasks (Zeng et al., 2024), where models perform similarly on agreement with or without metric-based guidance. This result suggests that evaluation capability is an internalized skill in modern LLMs, and external guidance provides only marginal benefits in relative assessment settings (Qin et al., 2024b). Furthermore, we find that reasoning models exhibit better performance under the Direct condition, indicating that imposing rigid structured metrics might constrain the models’ inherent reasoning processes, thereby limiting their evaluation potential. Furthermore, we validate the hypothesis from earlier: in single answer grading, the binary Rubric approach substantially outperforms the multi-point Likert scale, reinforcing that verifiable evaluation protocols yield more reliable judgments when there is a lack of relative information.
Table 4: Impact of observation forms on the performance of multimodal evaluators. The numbers in parentheses indicate the performance change relative to the setting with both code and image inputs.
Model Single Answer Grading Pairwise Comparison Image Only Code Only Both Image Only Code Only Both Claude-4-sonnet 54.43 ($\downarrow$ 1.07) 58.87 ($\uparrow$ 3.37) 55.50 57.34 ($\downarrow$ 8.72) 64.83 ($\downarrow$ 1.23) 66.06 GPT-4.1 53.98 ($\downarrow$ 4.89) 59.94 ($\uparrow$ 1.07) 58.87 57.34 ($\downarrow$ 8.72) 64.98 ($\downarrow$ 1.08) 66.06 GPT-4o 50.31 ($\downarrow$ 5.65) 56.57 ($\uparrow$ 0.61) 55.96 59.48 ($\downarrow$ 4.89) 63.46 ($\downarrow$ 0.91) 64.37 Gemini-2.5-pro 52.29 ($\uparrow$ 3.82) 48.93 ($\uparrow$ 0.46) 48.47 59.17 ($\downarrow$ 3.06) 60.55 ($\downarrow$ 1.68) 62.23 Qwen-2.5-VL-72B 50.00 ($\uparrow$ 1.07) 49.39 ($\uparrow$ 0.46) 48.93 57.19 ($\downarrow$ 6.42) 62.39 ($\downarrow$ 1.22) 63.61
Influence of Observation Modality
We next analyze how the input modality affects multimodal evaluators, comparing three observation forms: (1) Image Only, providing only the initial screenshot of the rendered web page; (2) Code Only, providing only the source code; and (3) Both, providing both code and screenshot. As shown in Table 4, code emerges as the most critical modality for evaluating web development tasks. Withholding code leads to a significantly larger performance drop than withholding screenshots, suggesting that while MLLMs can process visual input, their judgments are fundamentally anchored in the structured source code (Wang et al., 2025). In pairwise comparison, using both modalities yields the best results, indicating that visual context provides complementary signals that aid in refining relative judgments (Chen et al., 2025a).
4.3 Error Analysis
To systematically diagnose the failure modes of LLM-based evaluators, we conduct a fine-grained analysis along three key dimensions: inherent biases, limitations in understanding functional equivalence, and shortcomings in feasibility analysis.
Table 5: Positional bias in pairwise comparison. Preference for specific position, consistency and the absolute difference in agreement rate ($\Delta$ AR) between original and swapped orders are reported.
Model Consistency First Second $\Delta$ AR
Direct Likert Direct Likert Direct Likert Direct Likert
Claude-4-sonnet 89.6 87.9 5.7 3.7 4.7 8.4 0.9 1.38
GPT-4.1 83.3 85.2 0.9 3.0 15.8 11.8 0.8 0.46
DeepSeek-V3-0324 84.1 83.5 11.6 7.3 4.3 9.2 0.5 1.38
Inherent Biases in Judgment
Prior studies (Ye et al., 2025) have revealed that LLM-as-a-judge exhibits various forms of bias, such as positional and verbosity biases. We focus specifically on positional bias. Despite explicit instructions to ignore order and remain objective, models still exhibit a systematic preference for responses in a specific position, as shown in Table 5. One might hypothesize that this bias emerges primarily in ambiguous cases where the two options are of comparable quality. However, our analysis of the label distribution for instances with inconsistent predictions reveals that the proportion of ties is not higher than that of wins or losses. This finding suggests that positional bias is not merely an artifact of ambiguity but rather an inherent deficiency in the models, and instruction alone is insufficient to eliminate these deeply embedded inductive biases.
Figure 4: An example illustrating the failure to recognize functional equivalence. The webpage element for ‘Presentation’ serves the same purpose as the required ‘Demonstration’ rating.
Limitations in Understanding Functional Equivalence
A critical requirement for reliable evaluation is the ability to recognize functional equivalence—i.e., determining whether different implementations satisfy the same underlying requirement. We find that evaluators often fail in such judgments, adhering strictly to literal interpretation rather than intent. For example, as illustrated in Figure 4, when instructed explicitly, human evaluators correctly recognize an element implemented with alternative text, whereas both LLM and agentic evaluators incorrectly reject it. This reflects a fundamental gap in contextual and pragmatic reasoning, limiting their applicability in open-ended domains where diversity in implementation is common.
Table 6: Performance on the WebDevJudge-Unit for feasibility verification. We report precision (P), recall (R), F1-score, and accuracy (Acc).
Model Observe P R F1 Acc UI-TARS-1.5 Traj. 82.4 70.3 75.8 75.1 GPT-4.1 Code 72.1 90.0 80.1 75.1 DeepSeek-V3 Code 69.6 93.5 79.8 73.7
Shortcomings in Feasibility Analysis
We further investigate the evaluators’ ability to accurately verify task fulfillment. Existing benchmarks such as AgentBench (Lù et al., 2025) are too general to assess web-specific judgment. To address this, we construct WebDevJudge-Unit, a targeted dataset of 502 test cases. Each instance consists of web code, a specific verification task, an expected result, and a label indicating whether the task is feasible. We evaluate both LLM-based (code-only) and agent-based (interaction-driven) evaluators on this dataset. From the results in Table 6, we identify complementary weaknesses: (1) For LLM evaluators, these models achieve moderate recall but suffer from low precision. While they can often identify relevant code segments, they are unable to verify actual execution outcomes, leading to false positives when the code appears relevant but does not correctly implement the desired functionality. (2) For agentic evaluators, they exhibit higher precision but lower recall. They are effective when they successfully execute a test plan, but sometimes fail to complete tasks due to limitations in navigation or state interpretation. Consequently, they incorrectly label feasible tasks as infeasible due to their own operational failures rather than actual shortcomings in the web implementation. This divergence highlights a fundamental trade-off: static code analysis lacks execution grounding, while interactive agents are constrained by their own operational reliability. An ideal evaluator would combine the comprehensive coverage of code-aware reasoning with the grounded verification of interactive testing.
Our analysis reveals that the core limitation of LLM-as-a-judge lies in fundamental deficiencies in calibration capability. Lacking this calibration, LLMs struggle to map abstract quality dimensions onto discrete scores and verifiable rubrics. Improving judge performance will require addressing these core competency gaps rather than merely refining evaluation protocols.
5 Conclusion and Future Work
In this work, we introduce WebDevJudge, a comprehensive benchmark for evaluating LLM-as-a-judge in web development. Unlike previous benchmarks, WebDevJudge supports both static code analysis and interactive agent navigation with high-quality preference labels. Our experiments demonstrate that current LLM-as-a-judge approaches cannot effectively substitute human evaluation, and we identify core bottlenecks hindering their performance. Since our primary focus is on evaluation and analysis within the general LLM-as-a-judge domain, we did not specifically optimize the overall framework structure. We leave the exploration of sophisticated agentic workflows and complex multi-round evaluations to future work.
Ethics Statement
The ethical considerations of our work are discussed in the context of the following aspects: (1) Data collection and use. We use publicly available datasets and self-generated data for evaluation. We ensure that the data is only used for academic research purposes and no personal data is involved. We strictly follow the license terms and conditions. (2) LLMs API. We comply with the terms of service of the LLMs API providers strictly, maintaining fair use.
Reproducibility statement
We provide detailed descriptions of our methods, datasets, and evaluation metrics in the main text and appendix to ensure transparency and reproducibility. Our benchmark construction process for WebDevJudge  is detailed in Section 3 and Appendix A. The experimental setup, including the models, evaluation paradigms, and specific protocols, is described in Section 4 and Appendix B. We also provide a complete description of the WebDevJudge-Unit dataset in Appendix C. Code and scripts are available to replicate the empirical results.
Appendix A Details in WebDevJudge Construction
A.1 Data Collection and Filtering Pipeline
Our data collection process commences with the webdev-arena-preference-10k dataset (Vichare et al., 2025; Chiang et al., 2024), which contains 10,501 user queries, each paired with two web implementations and a user-provided preference label. To ensure the quality and suitability of this data for our benchmark, we implemented a rigorous two-stage filtering pipeline.
Query-based Filtering
The raw queries were first processed to address quality issues. We began by removing all verbatim duplicate queries to eliminate redundancy. Subsequently, we employed gemini-2.5-pro to screen for and exclude any queries containing harmful, offensive, or nonsensical content, based on a predefined set of safety and clarity instructions. We then exclude queries with minimal interaction requirements (score >> 8, max = 10) and queries with unclear intentions (score << 3, max = 5). The prompts used for the safety filter, interaction filter, and intention filter are presented below.
Prompt for Query-based Interaction Filtering Please analyze the user’s input to assess the level of detail regarding interaction requirements for web design. Focus on the functionality implied by the user’s input and the interactions required. Based on your evaluation, assign a score from 0 to 10: 0: No potential interaction requirements are mentioned, but the website may still contain interactions. 3: No interaction is mentioned explicitly, but interactions may still exist in the functionality. 5: One interaction is mentioned or implied, focusing on a single functionality or task. 8: More than one interaction is mentioned or implied, involving multiple functionalities or tasks. 10: All relevant interactions and functionalities are clearly mentioned in detail, requiring full user engagement with the site or app. After completing your analysis, output your score in the format: <<score:int>>. For example: <<score:7>>. Examples: User: Building a merge sort function Response: This is a programming task that doesn’t involve any web interactions or user-facing functionality. <<score:0>> User: a poker site Response: The term “poker site” implies interactions, but the input lacks specific details about those interactions. <<score:2>> User: Create a unique and incredibly beautiful design. The design is up to your discretion, but the main thing is to make it stand out! Write code that is at the senior level or even above it, as if you’ve been working in the frontend for 10 years. Make the design beautiful and pleasing to the eye! Response: The focus is on aesthetics and quality, with some implied user interaction related to the design’s appeal, but not much about functionality. <<score:3>> User: site for selling craft chocolate bars Response: This request implies basic interactions such as browsing and purchasing, but lacks specific details about the interactions. <<score:5>> User: Design a guided meditation player with progress tracking. Create a library view to browse and select different meditation sessions. Response: The request involves multiple interactions like browsing, selecting sessions, and tracking progress, but is not fully detailed. <<score:8>> User: Build a game of chess Response: A chess game involves numerous interactions such as moving pieces, taking turns, and tracking the game state, all requiring user engagement. <<score:9>> User: Make me a clone of WhatsApp Chat App Response: A clone of WhatsApp implies multiple interactions, including sending/receiving messages, media sharing, and managing chats, all of which require significant user engagement and functionality. <<score:9>>
Prompt for Query-based Safety Filtering You are given a user query. Classify whether it is safe or unsafe. Definitions: SAFE: The query is clear, relevant, and does not request or promote hazardous, violent, hateful, or sexually explicit content. UNSAFE: The query contains harmful, offensive, or nonsensical content. Output your decision in **valid JSON** with the following schema:
  "query": "<the original user query>",
  "label": "SAFE" or "UNSAFE",
  "category": "<one of: hazard, violence, sexual, hate, other>",
  "reason": "<brief explanation>"
Prompt for Query-based Intention Filtering Analyze the user’s input to evaluate the clarity of their query regarding website design. Assign a score from 1 to 5 based on the criteria below: Score 5: User provides a clear, detailed description of requirements, including specific features and design elements. Score 4: User’s intention is mostly clear with some details, but lacks comprehensive specifics. Score 3: User presents a general idea, but the request is vague and lacks essential information. Score 2: User’s input is unclear or contains irrelevant information, making their intention difficult to discern. Score 1: Input is nonsensical or purely code-related, reflecting no intention for a design request. Output: Return a JSON object containing the reasoning for the score and an expected result, formatted as follows:
        "reason": "A clone of WhatsApp Chat App",
        "score": "4"
Environment-based Filtering
Following the query-based filtering, we proceeded to validate the web implementations. We established a standardized Next.js environment for deployment. To manage dependencies, we identified and included the most common packages, thereby excluding implementations that required niche or incompatible libraries. Each successfully deployed implementation was then verified via a test request; those returning a non-200 status code were discarded as buggy. Finally, to handle instances with intrinsic runtime errors not captured by status codes (e.g., rendering a blank page), we captured an initial screenshot of each webpage. A multimodal model, Qwen2.5-VL-72B, was utilized to visually inspect these screenshots and filter out any pages exhibiting rendering failures.
This comprehensive filtering process yielded a final set of 1,713 high-quality instances for subsequent annotation. While the original dataset includes user-provided preference labels, we identified several factors that render them unreliable for rigorous evaluation. Primarily, the labels are susceptible to a high degree of subjectivity and variance in evaluation criteria among individual users. The single-pass nature of the crowdsourced annotations also lacks the verification necessary for robust benchmark data. Furthermore, the dataset includes a “tie (bothbad)” category, which introduces ambiguity. The subjective definition of “bad” can lead to inconsistent labeling, potentially masking instances where one implementation, though imperfect, is demonstrably superior to the other. To illustrate these inconsistencies, we provide examples of problematic cases from the original dataset in Table 7.
Table 7: Examples of problematic cases. In each example, the upper figure is referred to as model_a, while the lower one is model_b.
Query Implementation Analysis Build a website iden- tical to apple.com. [Uncaptioned image] Initial label: model_b Annotated label: model_a Problem type: subjective label bias. Reason: A thorough check of apple.com shows that A’s rest- oration is much more accurate. [Uncaptioned image] Build a 4×8 version of the game 2048. [Uncaptioned image] Initial label: tie (bothbad) Annotated label: model_b Problem type: misjudgment in “bothbad” ties. Reason: While both have clear usability flaws, B is still clearly superior to A in terms of its ali- gnment with the query. [Uncaptioned image]
A.2 Rubric Annotation Pipeline
The annotation was conducted by two of the authors, all of whom possess strong backgrounds in computer science and software engineering. Since we have two annotators, the agreement rate is calculated as the proportion of instances where their annotations are the same. The annotation requires annotators to interact with the deployed web implementations and provide a preference label via pairwise comparison. However, this direct approach yields low inter-annotator agreement, underscoring the high degree of subjectivity inherent in the task and the need for a standardized evaluation framework.
To address this, we introduced a rubric-guided annotation process. Manually creating detailed rubrics for each query is not only prohibitively time-consuming—with an estimated 20 minutes per rubric, excluding research time—but also intellectually demanding. It requires extensive background knowledge, including familiarity with established UI/UX design patterns for similar applications and, in many cases, specialized domain knowledge for tasks like scientific simulations. Recognizing these challenges, we leveraged a powerful large language model, gemini-2.5-pro, for automated generation. Based on each user query, the model produced a structured rubric tree organized along three core dimensions critical to web development quality:
  • Intention: The core requirements of the user query.
  • Static Quality: The assessment of static elements, including UI layout and UX design.
  • Dynamic Behavior: The evaluation of interactive features.
The prompt used to generate these rubric trees is provided below:
Prompt for Rubric Tree Building ## TASK DESCRIPTION
You are an expert software quality assurance (QA) analyst. Your task is to take a user query for a web development project and generate a structured, hierarchical rubric. This rubric will be used to evaluate a generated webpage in a verifiable, binary (implemented/not implemented) manner.
The output must be a single JSON object with ‘‘‘json and ‘‘‘ wrapped around it. The JSON object must have three top-level keys: ‘intention’, ‘static’, and ‘dynamic’. ### JSON Structure Rules: 1. Each node in the tree must be a dictionary with two keys: - ‘description’: A string describing the feature or goal. - ‘children’: A list of child nodes, or None if it is a leaf node. 2. The ‘intention’ section should capture the high-level purpose and core goals of the webpage. Descriptions should be concise and conceptual overviews of what the user wants to achieve. 3. The ‘static’ section must detail all the non-interactive, visible elements of the webpage. Break down components into their smallest logical parts. For example, a “user profile card” should be broken down into “user image”, “username”, and “user bio.” 4. The ‘dynamic’ section describes all the interactive functionalities of the page. - It must have exactly two children: one for “basic” interactions and one for “complex” interactions. - ‘basic’: These are simple, single-step user actions. Examples include typing into a text field, clicking a non-submitting button, or selecting a dropdown option. - ‘complex’: These are multi-step processes or actions that result in a significant change to the application’s state. Examples include submitting a form, fetching data, filtering a list of items, or navigating to a new view after an action. 5. Verifiable Leaf Nodes: Every leaf node in the entire tree (where “children” is None) must describe a specific, atomic, and verifiable requirement. The description should be a clear statement that can be evaluated as “implemented” or “not implemented”. ## Example: ### User Query:
{example_query}
### Generated Rubric Tree (Your Output):
‘‘‘json
{example_rubric_tree}
Now, analyze the following user query and generate the rubric tree in the specified JSON format. ### User Query:
{user_query}
Upon manual review, the LLM-generated rubrics exhibited both strengths and weaknesses. For queries that referenced real-world applications, the rubrics were often of high quality, sometimes surpassing human-authored versions in detail. However, for vague or overly general queries, which are common on crowdsourcing platforms, the generated rubrics sometimes included criteria that were either too specific or too broad, reflecting the ambiguity of the original request. We also provide the statistics of the rubric trees; the details are shown in Table 8. From the statistics, especially the average number of leaf nodes, we can see that the LLM-generated rubrics cover a wide range of criteria, which is a good sign for the evaluation of web development tasks.
Table 8: Statistics of the generated rubric trees.
Metric Intention Static Dynamic Whole
Average height 2.0 3.9 3.3 5.0
Average number of leaf nodes 3.6 15.6 10.6 29.9
To ensure consistent application of these rubrics, we established a set of clear annotation guidelines for the expert annotators, as detailed in the following:
Annotation Guidelines 1. All judgments must be based on actual user experience, not solely on visual appearance. 2. In general, the completeness of functionality takes precedence over aesthetics, unless the user query explicitly requests a focus on design. 3. If both web pages are well-implemented, aesthetic quality should be considered as a deciding factor. 4. If there is any important and discernible difference in quality between the two pages, a preference should be stated rather than defaulting to a tie. 5. When a decision is difficult, the rubric tree should serve as the definitive guide. Refer to the fulfillment of both leaf-level and root-level criteria. 6. When judging based on the rubric tree, consider functional equivalence rather than demanding literal, identical implementations.
This structured pipeline, combining LLM-generated rubrics with clear human oversight and guidelines, proved highly effective. The final inter-annotator agreement rate reached 89.7%, confirming the consistency and objectivity of our annotation process.
A.3 Data Statistics
In this section, we describe the categorization pipeline for WebDevJudge and present representative examples for each subcategory. We adapt topics from the original dataset (Vichare et al., 2025) to serve as our subcategories. Given that WebDevJudge contains only 654 instances, we did not use a topic modeling model for clustering. Instead, we generated subcategories by providing the topic name, a detailed description, and the user query to GPT-4o. These subcategories were then manually reviewed and consolidated into three main categories.
Appendix B Experimental Details
In this section, we show the details of the implementation of our experiments. Including the models and hyperparameters, the details of evaluation protocols and metrics, and the agentic workflow implementation.
B.1 Models And Hyperparameters
Our experiments utilize a combination of open-source and commercial models. Open-source models were deployed using vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024). Commercial models were accessed via their respective APIs, including Azure [^1] https://learn.microsoft.com/en-us/azure/cognitive-services/openai/reference and Vertex AI [^2] https://cloud.google.com/vertex-ai/generative-ai/docs/learn/overview. To ensure reproducibility, we set the temperature to 0.0 for all models, with the exception of claude-3.7-sonnet and claude-4-sonnet, for which the temperature was set to 1.0 to enable their “thinking” mode. For brevity, the models referred to as DeepSeek-V3 and DeepSeek-R1 in the main paper correspond to DeepSeek-V3-0324 and DeepSeek-R1-0528, respectively.
B.2 Evaluation Protocols and Metrics
Likert Scale
Motivated by the work of Bian et al. (2025), we designed a multi-level Likert scale for evaluating web development tasks. The scale is based on the international standard for software quality assessment (ISO/IEC/IEEE, 2022) and has been specifically adapted for our dataset. The details of the Likert scale are shown below:
Dimensions of the Likert Scale Evaluation Criteria
1. Functional Correctness and Completeness
- 1.1 **Core Functionality**: Evaluates if the primary features and requirements specified in the user query are implemented correctly and function as expected. - 1.2 **Content Accuracy and Completeness**: Assesses if all the required content (text, images, links, etc.) is present, accurate, and correctly placed as per the user’s query. - 1.3 **Boundary Conditions and Corner Cases**: Examines the solution’s behavior with unexpected or extreme user inputs. - 1.4 **Error Handling**: Evaluates the system’s ability to handle errors gracefully. This includes providing clear, user-friendly error messages and preventing application crashes due to invalid operations. 2. User Interface Quality - 2.1 **Visual Consistency and Cohesion**: Assesses the consistency of design elements such as color schemes, typography, spacing, and component styling throughout the webpage. - 2.2 **Layout, Structure, and Responsiveness**: Evaluates the overall layout and structural organization of the content. This also critically assesses the responsiveness of the design across different screen sizes (desktop, tablet, mobile). - 2.3 **Aesthetic Appeal**: Assesses the overall visual appeal of the webpage. This includes the effective use of color, typography, imagery, and whitespace to create an engaging and modern user interface. 3. Code Quality - 3.1 **Readability and Maintainability**: Assesses the clarity and organization of the code. This includes proper indentation, meaningful variable names, comments where necessary, and a logical file structure. - 3.2 **Modularity and Reusability**: Evaluates whether the code is broken down into logical, reusable components or functions, avoiding monolithic structures and code duplication. - 3.3 **Scalability and Efficiency**: Assesses the efficiency of the code, as well as the ability to scale the codebase for future enhancements or new features. 4. Interactivity: - 4.1 **Effectiveness**: Assesses the functionality and user experience of interactive elements like buttons, forms, menus, and sliders. This includes visual feedback on user actions (e.g., hover states, loading indicators). - 4.2 **Logical Correctness**: Evaluates whether the application state changes correctly in response to user interactions. - 4.3 **Accessibility**: Evaluates how easy and intuitive it is for a user to navigate the webpage and interact with its elements to achieve their goals.
In selecting these dimensions, we considered the typical workflow of human evaluators in web testing. Given that the tasks in our dataset are primarily front-end focused, and acknowledging the current limitations of generative models in producing efficient, full-stack solutions, we have concentrated on fundamental aspects of web development, omitting higher-level criteria such as backend performance and efficiency.
For single-implementation evaluation, the model assigns a score from 1 (lowest) to 5 (highest) to each sub-dimension. The final score is the sum of all sub-dimension scores. For pairwise comparison, both implementations are presented to the model simultaneously, which allows it to assign scores based on their relative merits. The final preference is determined by the score difference between the two implementations. A preference is declared for the higher-scoring implementation if the score difference exceeds a threshold of 2; otherwise, the outcome is considered a tie.
Rubric
For rubric-based evaluation, the process is guided by the LLM-generated rubric tree. The model is provided with the rubric, the user query, and the web implementation(s). To mitigate the inherent ambiguity in rubric-based assessments, we instruct the model to recognize functional equivalence. For instance, a requirement for a heading is considered met if a heading element is present, even if its text or styling differs from a literal interpretation of the rubric.
In the single-answer grading setting, the model assigns a binary label (pass or fail) to each leaf node of the rubric. We then compute a pass rate of the leaf node for each of the three primary dimensions: intention, static, and dynamic. The final score is the weighted sum of these pass rates. For pairwise comparison, the model evaluates the two implementations against each leaf node, determining which is superior or if they are tied. This yields a win rate for each implementation of the leaf nodes across the three dimensions. A final score for each is calculated as a weighted sum of these win rates. Preference is awarded to the implementation with the higher final score. In our implementation, all dimension weights are set to 1.
Direct
For direct evaluation, instead of providing any criteria, we instruct the model to directly output its preference (i.e., Wa, Wb, or tie). The prompt we used is shown below.
Prompt for direct comparison You are tasked with comparing two React code snippets (Model A and Model B) based on the user’s query. The input will contain: - User Query: The question or concern the user has. - Answer from Model A: The output of Model A. - Answer from Model B: The output of Model B. Your output should be in the following JSON format: ‘‘‘ { “reason”: “Detailed explanation for why one model is better, based on the user’s query”, “winner”: “‘model_a’, ‘model_b’, or ‘tie”’ } ‘‘‘ - reason: A detailed explanation of why one model’s answer is better than the other, based on the user’s query. Focus on the quality of the responses and how well they address the user’s concerns. - winner: Select the winner from the following options: - “model_a”: If Model A is better. - “model_b”: If Model B is better. - “tie”: If both models are equally good or both provide unsatisfactory answers. Ensure that you thoroughly evaluate the responses before selecting the winner and providing the reason.
B.3 Agentic Workflow Implementation
This section details our agentic workflow implementation. To ensure comparability with our other evaluation methods, the workflow’s planner adopts the LLM-generated rubric tree to guide the executor’s verification actions. The executor is UI-TARS-1.5 (Seed, 2025), one of the state-of-the-art end-to-end GUI agents operating on a ReAct-style paradigm (Yao et al., 2023), where it first generates a thought and then a corresponding action. These actions are converted into executable pyautogui [^3] https://pyautogui.readthedocs.io/en/latest/ code. We utilize the official UI-TARS-1.5 action space and have designed specific prompting strategies for each rubric dimension. For the Static dimension, the rubric tree is converted into a list of elements; the agent navigates the page to find these elements and returns a list of those present. For the Dynamic and Intention dimensions, each rubric leaf node becomes a task for the agent to complete. Upon reaching the maximum number of steps, as detailed in Table 9, the agent provides a conclusion on the task’s outcome. Finally, the summarizer calculates a pass rate for each dimension based on the agent’s findings, and the final score is the weighted sum of these pass rates, mirroring the single-implementation rubric evaluation.
Table 9: Details of the agent settings.
Dimension Input Format Output Format Max Step
Static List of elements to check List of found elements 6
Dynamic Task description Task feasibility and conclusion 5 for basic, 15 for complex
Intention Intention description Feasibility conclusion 15
Action Space of UI-TARS 1.5 click(point=‘<point>x1 y1</point>’)
left_double(point=‘<point>x1 y1</point>’)
right_single(point=‘<point>x1 y1</point>’)
drag(start_point=‘<point>x1 y1</point>’, end_point=‘<point>x2 y2</point>’)
hotkey(key=‘ctrl c’) # Split keys with a space and use lowercase. Also, do not use more than 3 keys in one hotkey action.
[…]
scroll(point=‘<point>x1 y1</point>’, direction=‘down or up or right or left’) # Show more information on the ‘direction’ side.
wait() # Sleep for 5s and take a screenshot to check for any changes.
[…]
Table 10: Statistics of the WebDevJudge-Unit dataset.
Error Type Proportion Example Description Error Reason Non-funtional element 35.9 Task: On a Question Review page, modify the text content of a displayed question within its editable text field. Unable to edit question content. Expected Result: The displayed question text updates to reflect the new input. Missing element 30.9 Task: Click the ‘Market’ tab or navigation item. Unable to find the Market interface. Expected Result: The ‘Market’ view is displayed, showing a list of cryptocurrencies. Prerequisite not met 15.2 Task: Select an answer for a multiple-choice question by clicking on one of the options. Unable to import the quiz correctly. Expected Result: The chosen answer option is visually highlighted. Loading issue 9.0 Task: Click a file name in the File Explorer sidebar. Unable to load file content. Expected Result: The corresponding file’s mock content is shown in an editor tab. Unreasonable outcome 5.4 Task: Scroll the chat display area to view older messages when content overflows. N/A Expected Result: The user is logged out. Ambiguous input 2.2 Task: Hover over a data point or segment in a Chart. Chart is not required in the webpage. Expected Result: A tooltip containing mock data relevant to the hovered chart element is displayed. Overly detailed task 0.9 Task: From a list of activities on the ‘Activities Screen’, click on a specific activity’s name or image. Unable to get the expected result based on the operation. Expected Result: The ‘Detail View’ for the selected activity is displayed, showing its full description. Missing animation 0.5 Task: Move the mouse cursor across the main 3D animated background area. The background 3D animation is not responsive. Expected Result: The 3D background animation dynamically responds to the cursor’s position or movement.
Appendix C WebDevJudge-Unit Dataset
In this section, we introduce WebDevJudge-Unit, a task-level dataset created to assess the capability of evaluators to verify task feasibility.
C.1 Dataset Construction
To construct the WebDevJudge-Unit dataset, we began by randomly sampling 105 queries from WebDevJudge. For each query, we prompted gemini-2.5-pro to generate up to five corresponding verification tasks, each with an expected result. We then generated the necessary HTML code for each query to facilitate easy deployment. Following deployment, each task was meticulously annotated for feasibility (true/false) by interacting with the live web application. For tasks identified as infeasible, we further annotated the specific error type and provided a detailed reason for the failure.
C.2 Dataset Statistics and Examples
The resulting dataset comprises 502 tasks derived from 105 unique queries. The feasibility labels are distributed as follows: 279 tasks are marked as feasible, and 223 as infeasible. Table 10 presents a detailed breakdown of the error types for the infeasible tasks, along with illustrative examples for each category.
Figure 5: The consistency between model predictions under pairwise comparison.
Figure 6: The consistency between model prediction under single answer grading.
Appendix D Additional Results
This section provides a more detailed analysis of LLM-as-a-judge performance on web development tasks, supported by supplementary experimental results.
D.1 Prediction Distribution Consistency
To understand the nature of model failures, we analyze the consistency of predictions across different evaluators. This investigation seeks to determine whether different models tend to err on the same set of instances—suggesting certain examples are inherently challenging—or if their errors are largely independent and model-specific. The results of this label consistency analysis are presented in Figure 5 and Figure 6.
Our analysis of prediction consistency across different models reveals a significant disparity between the two evaluation paradigms. As illustrated in Figure 5 and Figure 6, the inter-model agreement under the pairwise comparison paradigm is substantially higher than that under single-answer grading. In pairwise comparisons, the consistency rates between different evaluators generally exceed 75%, with many model pairs even surpassing 80%. In stark contrast, under single-answer grading, inter-model consistency drops significantly, typically hovering between 50% and 65%.
This discrepancy corroborates our core finding from the main text: pairwise comparison is a more stable and reliable paradigm for complex, open-ended tasks like web development. Relative judgment in pairwise comparison constrains the evaluation scope, compelling the model to focus on discriminative features between two candidates, which is a cognitively less demanding task than absolute scoring. Conversely, single-answer grading requires the model to possess a well-calibrated and consistent internal standard of quality, a standard that varies greatly across different models. Consequently, the single-answer grading paradigm exposes the current models’ deficiencies in calibration, leading to inconsistent and less reliable evaluation outcomes.
Furthermore, the high consistency in pairwise comparison suggests that while different models may lack a shared understanding of ’absolute quality,’ their internal mechanisms for making ‘relative preference’ judgments are more aligned. This reinforces our view that leveraging relative judgments is a more robust approach for automated evaluation in domains characterized by nuance and ambiguity, where clear-cut right or wrong answers are scarce.
D.2 Mitigating the Impact of Positional Bias
To mitigate the impact of positional bias, we employ a widely-used debiasing technique from prior works (Zeng et al., 2024; Tan et al., 2025). This method involves evaluating each pair of implementations twice, with their positions swapped in the second evaluation. A preference is considered final only if the model’s choice remains consistent across both orderings. If the choice is inconsistent, the outcome is recorded as a tie. We use this approach to investigate whether mitigating positional bias improves overall model performance. The results are shown in Table 11.
Table 11: Agreement rate (%) with and without mitigating the positional bias.
Model w/ mitigating w/o mitigating
Direct Likert Direct Likert
Claude-4-sonnet 65.0 65.1 66.1 66.1
GPT-4.1 64.8 65.1 64.2 66.1
DeepSeekV3-0324 60.8 62.2 62.5 63.6
A natural question arises as to why we did not employ the swap-based debiasing technique in our main experiments. Our decision to rely on single-pass evaluation stems from two primary considerations. First, our core objective is to characterize the raw, unfiltered behavior of LLMs-as-judges to understand their inherent biases. Applying debiasing techniques from the outset would mask or average out these effects, obscuring the very phenomena we seek to analyze. The results in Table 11 indicate that the overall performance difference before and after debiasing is not substantial in our setup. This finding reinforces the importance of studying the biases directly, as their presence is not always evident from aggregate performance metrics alone. Second, the single-pass evaluation mirrors many practical application scenarios where, for reasons of cost and efficiency, models are queried only once per pair. Therefore, our main experimental results offer a more realistic and cautionary benchmark for practitioners, highlighting the potential pitfalls of naively deploying these models without safeguards. Our methodology thus deliberately separates the diagnosis of inherent model behaviors from the evaluation of mitigation strategies.
D.3 Evaluation Results on Non-Tie Cases
Given the brittleness of language models when handling ambiguous tie conditions, we also present results for non-tie cases, where a clear preference was established. As presented in Table 12, excluding tie cases results in a substantial improvement in agreement rates for all evaluators, indicating that ambiguous comparisons are a primary source of error. More importantly, these results reinforce the core conclusions drawn from our main experiments. The pairwise comparison paradigm consistently and significantly outperforms single-answer grading, highlighting its robustness for capturing relative quality. Furthermore, the patterns of different guidance mechanisms remain consistent: in the pairwise setting, both Likert scale and Rubric-based guidance yield comparable performance, suggesting that the relative judgment itself is the dominant factor. In contrast, for single-answer grading, the Rubric-based approach is demonstrably superior to the Likert scale. This reinforces our finding that structured, binary assessments provide a more reliable signal for absolute evaluation than multi-point scales, which require a level of calibration that models currently lack.
Table 12: Agreement Rate (%) (without tie) of different evaluators under different evaluation paradigms.
Model/Method Digital Design Game & App Web & Special Average Single Pair Single Pair Single Pair Single Pair Likert  GPT-4.1 65.73 80.34 71.27 82.87 72.61 85.35 69.77 82.75  GPT-4o 58.99 78.65 64.09 80.11 70.70 80.89 64.34 79.84  Qwen-2.5-VL-72B-Inst. 50.00 74.72 51.93 76.24 59.87 78.98 53.68 76.55  Gemini-2.5-flash-lite 56.74 69.10 56.91 69.61 57.32 73.89 56.98 70.74 DeepSeek-V3-0324 67.42 78.09 64.09 75.14 71.34 83.44 67.44 78.68 Kimi-K2-Inst. 69.66 78.65 66.30 73.48 63.69 80.25 66.67 77.33 Qwen3-235B-A22B-Inst. 62.36 75.28 67.40 77.35 67.52 75.80 65.70 76.16 Qwen3-30B-A3B-Inst. 65.73 76.97 69.06 73.48 67.52 81.53 67.44 77.13 Qwen2.5-32B-Inst. 55.06 71.91 58.56 77.35 59.24 77.71 57.56 75.58 Gemma-3-27B-it 51.12 69.66 57.46 64.09 49.04 75.16 52.71 69.38 Qwen2.5-14B-Inst. 51.69 59.55 50.28 56.35 43.31 59.87 48.64 58.53  Claude-4-Sonnet 61.24 80.90 69.06 82.32 69.43 80.25 66.47 81.20  Claude-3.7-Sonnet 75.28 79.78 69.61 75.69 69.43 85.35 71.51 80.04  Gemini-2.5-pro 59.55 75.28 58.56 79.01 50.32 73.25 56.40 75.97 GLM-4.5 67.42 79.78 63.54 76.80 64.33 78.98 65.12 78.49 DeepSeek-R1-0528 57.87 76.97 65.75 76.80 58.60 76.43 60.85 76.74 Qwen3-30B-A3B-Think. 62.36 71.35 61.88 72.38 61.78 75.80 62.02 73.06 Rubric  GPT-4.1 78.09 82.02 70.72 73.48 70.70 87.26 73.26 80.62  GPT-4o 75.84 78.09 71.27 75.69 73.25 83.44 73.45 78.88  Qwen-2.5-VL-72B-Inst. 70.79 76.40 71.27 77.9 73.25 80.89 71.71 78.29 DeepSeek-V3-0324 78.09 79.21 65.75 79.01 78.34 87.26 73.84 81.59 Kimi-K2-Inst. 72.47 82.02 67.40 74.59 75.16 87.26 71.51 81.01 Qwen3-235B-A22B-Inst. 74.16 76.97 65.75 66.3 71.97 75.16 70.54 72.67 DeepSeek-R1-0528 78.09 79.21 64.09 70.17 73.25 77.07 71.71 75.39 GLM-4.5 76.97 80.34 66.30 70.72 77.71 79.62 73.45 76.74 Agentic workflow 67.98 - 71.27 - 75.80 - 71.51 -
Appendix E The use of large language models
We use a large language model (LLM) as a general-purpose writing assistant. The primary use of the LLM is to refine and improve the clarity and flow of the text. Specifically, we provided the LLM with our pre-existing draft text and a clear outline of our ideas. We instruct the model to adhere strictly to the provided content, without adding or removing any information. The LLM’s role is limited to enhancing the logical coherence and fluency of the language, ensuring the final text is more polished and easier to read.